IN A YOUNG DOULBLE SLIT EXPERIMENT THE INTENSITY AT CENTRAL MAXIMUM IS I. THE INTENSITY AT A DISTANCE B/4 AT CENTRAL MAXIMUM IS ( B IS THE FRINGE WIDTH)

Grade:Upto college level

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Arun Dash,

Ans: B=λD/d................(1)

Now at a distance B/4 from the central maxima the angle made by the light ray is θ, then

tanθ=B/4D≈θ≈sinθ (as θ is very small).......................(2)

Now path difference is d sinθ=d B/4D and hence phase difference is (2∏/λ)d (λD/d)/4D=∏/2

Now intensity at this point is=Icos²(∏/2)=I/2

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Arun Dash !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur